OpenStudy (anonymous):

Find dy/dx. y=5^(ln(x^2)+3)

OpenStudy (abb0t):

\[a^x \ln(a)\]

OpenStudy (anonymous):

thank you!:)

OpenStudy (abb0t):

fa sho! GL!

OpenStudy (raden):

lol, the clue incomplete... use chain rule : y=x^2 ---> y'=2x v=ln(u) ---> v'=1/u w=5^v ---> w'=5^v * ln(5)

OpenStudy (anonymous):

ok, i got this.. y'= 5^(ln((x^2)+3)) *ln 5*((2x)/((x^2)+3)) ..is it right?

OpenStudy (raden):

for the power of 5, is {ln(x^2) + 3} or ln(x^2+3)

OpenStudy (anonymous):

i did include that, im just not sure if i did it right?, do i have to put (1/5) for ln5?

OpenStudy (raden):

before verify, i want to asking u is ur question